# Understanding and Imitating Human-Robot Motion with Restricted Visual Fields

### v2

@@ Abstract
When working around humans, it is important to model their perception limitations in order to predict their behavior more accurately. In this work, we consider agents with a limited field of view, viewing range, and ability to miss objects within the viewing range (e.g., transparency). By considering the observation model independently from the motion policy, we can better predict the agent’s behavior by considering these limitations and approximating them. We perform a user study where human operators navigate a cluttered scene while scanning the region for obstacles with a limited field of view and range. Using imitation learning, we show that a robot can adopt a human’s strategy for observing an environment with limitations on observation and navigate with minimal collision with dynamic and static obstacles. We also show that this learned model helps it successfully navigate a physical hardware vehicle in real-time. Code available at https://github.com/labicon/HRMotion-RestrictedView.
@@ I INTRODUCTION
As robots become more ubiquitous, they need to be able to operate in environments and use tools designed for humans as well as predict the behavior of humans in their proximity. In order to predict the motion of a human, the capabilities and limitations of a human must be understood by the robot. One aspect is localized attention, where a human’s limited field of view and head orientation can be used as direct indicators of what the human is currently aware of from visual perception. In this work, we argue that accurately modeling, predicting, and imitating human behavior requires us to account for the inherent limitations in human perception and observation abilities.
Consider, for example, a humanoid robot whose motion-action and observation are distinct. To navigate an unknown environment, the robot must turn its head to be aware of static and dynamic obstacles in the environment [1]. Studies have been performed that observe the motion of human athletes while they play games such as basketball. In these tasks, it has been shown that the player’s attention can be inferred by approximating their gaze and can be good indicators of expected behavior [2, 3]. Even more generally, human intentions can be inferred from their gaze direction given their limited field of view [4, 5, 6, 7]. These studies suggest that a human’s field of view and point of view play a key role in understanding human intentions and predicting their actions.
@@
Figure 1: Illustrative Example. We present a robotic agent whose motion policy and observation ability are decoupled. The robot may move its base (orientation $\Phi$) and independently move its camera ($\Psi$). The observation of the camera has a limited range $r_{obs}$ and field of view $\Theta_{\text{obs}}$ and the likelihood of observing an object ($O_{i})$ within the field of view. The goal is to navigate to the goal without collision, and this work postulates that the correct modeling of both the policy and observation model is necessary to predict the motion behavior of the agent accurately.
@@
Figure 2: We employ the diffusion policy learned from human data on a car. The car successfully navigates the environment with obstacles. Like humans, the car changes its observation direction to look around, detect obstacles, and avoid them while moving toward the goal in real-time.
There has been extensive literature on learning from human demonstrations and imitating them. An agent’s policy can be approximated and replicated using methods from Learning from Demonstration [8, 9, 10]. By observing a demonstrator, the agent’s reward function can be inferred from their actions (Inverse Reinforcement Learning - IRL) [11]. Alternatively, one can directly learn and predict the actions of the agent using Behavior Cloning Strategies (BC) [12, 13, 14, 15] or hybrid approaches where a discriminator must distinguish between artificial actions from a generator and actual demonstration actions, and the generator is capable of sufficiently confusing the discriminator is a trained agent [16, 17]. When it comes to BC methods, Diffusion models have seen a recent rise in their application to learning from the demonstration as actions are diffused from noise conditioned on observations and states and accommodate multi-modality  [18, 19, 20, 21, 22]. While these models have proven to be successful in modeling human choices and intentions in many settings, we argue that to better model, predict, and imitate agents such as humans, we need to account for the perception capabilities and observation space of the agents as well. Leveraging these models, we argue that robots that can approximate the observation of another (human) agent, including the field of view, head/gaze/camera orientation, and the likelihood of observing an object, in addition to the motion policy, to better predict the actions of that agent since the agents’ motion policy will be conditioned on their observations.
@@
Our key idea is that to better model, predict, and learn agents’ motion and policy, it is necessary to correctly model both the observation model of the agent and their policy. There are two key components that affect an agent’s perception of an environment, which ultimately affect an agent’s actions: (a) the field of view and range of the agent’s visual observation, which determines where the agent is looking and what it can potentially perceive, and (b) the likelihood of successfully detecting objects within that observation region by the agent. In this paper, we examine the role of these two components and show that we can infer and estimate these parameters using demonstration trajectories of an agent. We further show that by reasoning about these parameters and accounting for them, one can generate trajectories that closely mimic the demonstrated trajectories.
@@
Our contributions are as follows: to better understand the relationship between an agent’s observation strategy and motion policy, we create a robotic agent with a limited field of view, view range, and likelihood of observing an obstacle and use the agent’s motion to infer these quantities. We consider three distinct scenarios, where each scenario involves one unknown: the parameters of the observation space, the probability of detecting obstacles within the agent’s field of view, or the agent’s motion policy, which depends on its observations. In all scenarios, we show that by accounting for the agent’s inherent observation limitations, we can more accurately model and predict its behavior. More specifically, we show that to infer the parameters of the observation space; one can use the cross-entropy [23] method to estimate the observation parameters accurately. We show that using Bayesian optimization [24], the probability of object detection within an agent’s field of view can be easily estimated. Finally, for inferring the policy of the agent, we show that the agent’s policy can be accurately imitated using diffusion models [19]. We illustrate this by training a diffusion policy on human demonstrations of navigating an environment under a restricted field of view. We demonstrate our ability to predict and replicate human behavior for both observation and motion in the presence of static and dynamic obstacles. Finally, we deploy the imitated policy to physical hardware, showcasing a robot navigating a cluttered environment and exhibiting human-like behavior by dynamically rotating its field of view to detect obstacles.
@@ II-A Problem Formulation
Consider an agent navigating in an environment with obstacles. The obstacles could be either stationary or moving. The agent operates under partial observability, as it can only perceive a limited portion of the environment at any time. This naturally aligns with a Partially Observable Markov Decision Process (POMDP) framework, where the agent maintains a belief over the true state of the environment based on incomplete observations.
We use $x_{t}\in\mathbb{R}^{n}$ to denote the state of the agent at time $t\in\mathbb{N}\cup\{0\}$, where is =$\mathbb{N}=\{1,2,3,\ldots\}$ is the set of natural numbers. Note that here, $n$ represents the agent’s state space dimension. Let the control action of the agent at time $t$ be denoted by $u_{t}\in\mathbb{R}^{m}$ where $m$ is the dimension of the control space of the agent. We consider discrete deterministic dynamics for the agent denoted through a function $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$,
@@
We assume the robot’s dynamics are known and that the agent has limited observation capabilities. The agent’s observation at time $t$ is denoted via $z_{t}\in\mathbb{R}^{o}$ where $o$ is the dimension of the observation space of the agent. We assume that the agent has an observation space that determines the area of the environment the agent is looking towards. Furthermore, we assume that the agent has an inherent probability of observing something present in its observation area. With the introduction of this probability, we try to entail the scenarios when humans sometimes miss objects even when they are looking in their direction. The combined observation space and probability of observation determine the agent’s observation at each time, given the agent’s current state. Let the observation space parameter that determines the agent’s observation area be denoted by $\theta\in\Theta$ where $\Theta$ is the space of observation space parameters. Also, let the probability of observing a particular observation be denoted by $p_{obs}\in[0,1]$. These components define the agent’s POMDP observation model, represented as
@@
At each time $t$, based on the observation $z_{t}$, the agent takes a control action $u_{t}$. The probability of taking an action at each time $t$ by the agent is captured through a stochastic agent policy, which we denote via $\pi:\mathbb{R}^{m}\times\mathbb{R}^{o}\rightarrow[0,1]$,
@@
Running Example: To illustrate this idea, consider an agent with discrete dynamics (Eqn. (1)) navigating through an environment with obstacles towards its desired goal location(Fig. 1). The agent visualizes the scene with a visual sensor parameterized by the range of vision ($r_{obs}$), and the field of view ($\theta_{obs}$). Obstacles in the field of view are not guaranteed to be seen (due to transparency and lack of visual features), which is captured by the detection probability $p_{obs}$. The motion of the agent is governed by the policy $\pi(\cdot;\theta_{\pi})$, which controls a) the motion of the robot in the environment, and b) the orientation of the visual sensor denoted by $\psi_{obs}$. With these constraints, the agent follows a POMDP-based navigation strategy, making decisions under uncertainty to minimize collisions while reaching the goal.
Given this example, we look at three key problems: 1) estimating the agent’s observation space parameters, 2) calculating the probability of detecting an obstacle within the field of view, and 3) estimating the agent’s motion policy for both displacement and sensor motion.
It should be noted that while an end-to-end learning approach could, in theory, capture the complexities of human-like navigation, our modular approach offers several advantages in terms of interpretability, generalization, and robustness. By explicitly modeling the agent’s observation space, detection probability, and motion policy separately, we gain deeper insights into the agent’s decision-making process and can analyze the influence of each component. This structure also enhances generalization, as the learned components can be adapted to different environments without requiring extensive retraining. Lastly, this decomposition allows for improved robustness, as failures in one module (e.g., imperfect perception) can be analyzed and mitigated without affecting the entire system, which is often challenging in end-to-end learning where errors can propagate unpredictably.
@@ II-B Estimating the Observation Space
During the estimation of the observation parameters $\theta$, we assume that the probability of observing an obstacle $p_{obs}$ within an observation space and the agent’s policy parameters $\theta_{\pi}$ are known.
@@
[…]
Remark 1
It should be noted that the integration sign in (II-B) will turn into a summation if the observation space of the agent is discrete.
From now on, we omit denoting that integration is over $z_{t}$ for the sake of brevity. From (II-B), we can compute the probability of the whole dataset $\mathcal{D}$ as $P(\mathcal{D};\theta)=\sum_{i}^{N}P(\tau_{i};\theta).$ When dealing with probabilities, we typically prefer to maximize the log-likelihood of probabilities of all individual trajectories. Therefore, the problem of estimating $\theta$ can equivalently written as
@@
Solving (5) requires taking the gradient of the cost function with respect to $\theta$, which would result in having to take the gradient of policy $P(u_{t}|z_{t})$ with respect to $\theta$ where $z_{t}$ depends on $\theta$. This is difficult because of the probabilistic dependencies of variables on each other. Therefore, we resort to black-box optimization techniques such as Bayesian optimization [24] and cross-entropy method [23], which find the optimal solutions based on evaluations of the function without requiring us to compute the gradient of the objective.
Bayesian optimization is a probabilistic model-based optimization technique used to optimize expensive objective functions to evaluate efficiently. In particular, we employ an open-source Python toolbox for Bayesian optimization [25] to solve (5). The Cross-Entropy Method is another optimization technique that iteratively samples random parameters, evaluates their performance, and refines the parameter distribution toward the optimal region [23].
@@ II-C Estimating the Probability of Observation
Under this assumption, we can compute the probability of a trajectory given a value of $p_{obs}$. Estimating $p_{obs}$ can be cast as:
@@
The problem in (6) is similar to that of (5). Therefore, we can use the same black-box function optimization techniques, such as Bayesian optimization or cross-entropy methods.
@@ II-D Estimating the Policy of the Agent
Under this assumption, the problem of estimating $\theta_{\pi}$ can be treated as a behavior cloning problem, i.e., we aim to learn the policy $\pi(\cdot;\theta_{\pi})$ parameterized by $\theta_{\pi}$ such that it imitates the provided data set $\mathcal{D}$ in the best possible way. If the policy is a neural network, then $\theta_{\pi}$ can be considered a neural network parameter.
State of the art in imitation learning is diffusion policies [20]. These methods model policy generation as a diffusion process, where an initial noisy signal or random trajectory is progressively refined toward optimal actions through iterative steps. As such, we adopt diffusion policies for mimicking the agent’s policy. We formulate policies such as Denoising Diffusion Probabilistic Models (DDPMs) conditioned on the state and observation of the agent. The detailed implementation for this method can be found in [20].
@@ III Numerical Simulations
In this section, we present numerical simulations to showcase the importance of considering agents’ observation and perception capabilities when either learning from them or modeling their behavior. In this section, we consider synthetic settings where we have access to the ground-truth parameters which enable a comprehensive comparison and evaluation of our method. Then, in the next section, we consider human data and perform hardware experiments that leverage actual human data. We consider a unicycle agent with discrete dynamics:
@@
Figure 3: Schematic for the unicycle navigation scenario with partial observations. The bright orange circle represents the unicycle. The gray circles represent the environmental obstacles, and the green represents the goal. The light orange cone represents the observation space of the agent, which is parameterized by the radius of the cone ($r_{obs}$) and the angle of the cone ($\theta_{obs}$). Each obstacle lights up red when the agent observes the obstacle.
@@
First, we consider the problem of estimating observation space parameters of the unicycle agent, $\theta=(r_{obs},\theta_{obs})$. We consider four different scenarios with different $(r_{obs},\theta_{obs})$ values and generate four datasets 500 trajectories each. During each trajectory generation, we randomly choose the location of the obstacles. We assume that the agent follows a known deterministic policy that is attractor-repulsor in nature, i.e., a policy that guides the agent to move toward the goal while avoiding obstacles. The visualization of one of the such trajectories is provided in Figure 3. We use the cross-entropy method to solve (5) and estimate the actual parameters $(r_{obs},\theta_{obs})$. For each set of parameters, we provide the ground truth values and the estimated values in Table I. As it can be seen, our method can closely estimate the actual parameter values.
@@
TABLE I: Actual and estimated values of the observation space parameters
We further examine the trajectories that these estimated estimated $\theta^{*}=(r^{*}_{obs},\theta^{*}_{obs})$ induce to examine how close they are to the demonstration data. We compute the Fréchet distance between the actual and predicted trajectories to investigate their similarity. The Fréchet distance is a good metric for comparing curves because it captures the overall shape and continuity by considering the closest possible continuous matching between the two curves.
@@
TABLE II: Average normalized Fréchet distance between 500 actual and predicted trajectories for scenarios in Table I
The Fréchet distances between actual and predicted trajectories for all scenarios are provided in Table II. As can be seen, for all four datasets, the estimated parameters are within about $6\%$ of the interval to the original trajectories. These results confirm that by reasoning about the observation space and perception capabilities of the agent, mimic the agent behavior and generate agent trajectories that are close to the demonstrated trajectories.
Next, we estimate the probability of observing the obstacles, $p_{obs}$. For this method, we consider the same policy as before, and we further assume access to the actual observation space parameters $(r_{obs},\theta_{obs})$. Same as before, we consider four different scenarios which correspond to four different values of $p_{obs}$ to generate four datasets of 500 trajectories each. Same as before, we randomly choose the locations of obstacles during each trajectory. We employ the Bayesian optimization method to solve (6) and estimate the $p_{obs}$ values. The actual values of $p_{obs}$ and their corresponding estimated values are presented in Table III. Our method can predict the $p_{obs}$ values within about $10\%$ of the actual values at the least.
@@
TABLE III: Actual and estimated values of observation probability, $p_{obs}$
@@
TABLE IV: Average normalized Fréchet distance between 500 actual and predicted trajectories for scenarios in Table III
Finally, we focus on estimating the agent’s policy. We generate a dataset of 500 trajectories with different obstacle configurations for all trajectories using the aforementioned attractor-repulsor policy with fixed values of $r_{obs},\theta_{obs}$, and $p_{obs}$. We aim to learn a policy that best imitates the dataset. For this purpose, we employ Diffusion policies where we learn an action diffusion policy that takes action based on observation and the agent’s current state. It should be noted that we do not have different scenarios for this part as we aim to learn only one policy. However, this can easily be extended to different policies. Using the learned diffusion policy, we induce a dataset of predicted trajectories. As before, we compute Fréchet distances between the actual and predicted trajectories. The average normalized Fréchet distance is $3.47\pm 9.31$, confirming that the resulting trajectories resemble demonstration trajectories.
To further assess the effectiveness and robustness of our modular approach, we evaluate the generalization capability of the learned diffusion policy. Specifically, we train our diffusion policy on a set of observation space parameters and test its performance on previously unseen observation parameters. We then compare the trajectories generated by the learned policy with those produced by the original policy. The results, summarized in Table V, show the average normalized Fréchet distances between the predicted and original trajectories across different observation parameters. Despite variations in the observation space, the learned policy successfully generalizes to new settings, producing trajectories that closely resemble those from the original policy. This demonstrates the adaptability of our modular approach, reinforcing its ability to capture underlying behavioral patterns while maintaining robust performance across diverse conditions.
@@
TABLE V: Average normalized Fréchet distance between 100 different trajectories generated by trained diffusion model against the original policy for different observation parameter values.

### v3

@@ Abstract
When working around other agents such as humans, it is important to model their perception capabilities to predict and make sense of their behavior. In this work, we consider agents whose perception capabilities are determined by their limited field of view, viewing range, and the potential to miss objects within their viewing range. By considering the perception capabilities and observation model of agents independently from their motion policy, we show that we can better predict the agents’ behavior; i.e., by reasoning about the perception capabilities of other agents, one can better make sense of their actions. We perform a user study where human operators navigate a cluttered scene while scanning the region for obstacles with a limited field of view and range. We show that by reasoning about the limited observation space of humans, a robot can better learn a human’s strategy for navigating an environment and navigate with minimal collision with dynamic and static obstacles. We also show that this learned model helps it successfully navigate a physical hardware vehicle in real-time. Code available at https://github.com/labicon/HRMotion-RestrictedView.
@@ I INTRODUCTION
As robots become more ubiquitous, they need to be able to operate in environments that are designed for humans as well as predict the behavior of humans and other agents in their proximity. In order to predict the motion of agents such as humans, the capabilities and limitations of a human must be understood by the robot. In this work, we argue that accurately modeling, predicting, and imitating human behavior requires us to account for the inherent limitations in human perception and observation capabilities.
Evidence from cognitive science and behavioral studies supports this view. In sports, studies show that human gaze and field of view are strong indicators of attention and intended actions [1, 2]. Even more generally, human intentions can be inferred from their gaze direction given their limited field of view [3, 4, 5, 6]. These studies suggest that a human’s field of view and point of view play a key role in understanding human intentions and predicting their actions.
@@
Figure 1: Illustrative Example. We present a robotic agent whose motion policy and observation ability are decoupled. The robot may move its base (orientation $\Phi$) and independently move its camera ($\Psi$). The observation of the camera has a limited range $r_{obs}$ and field of view $\Theta_{\text{obs}}$, and the likelihood of observing an object ($O_{i})$ within the field of view. Our goal is to accurately predict the motion of the agent navigating through the environment by observing its motion. In this work, we propose that the correct modeling of both the policy and observation model is necessary to predict the motion behavior of the agent accurately.
Building on these findings, imitation learning seeks to replicate agent behavior, yet doing so effectively requires accounting for the demonstrator’s perceptual limitations. Current literature learns from demonstrations [7, 8, 9] primarily through three approaches: inferring reward functions (Inverse Reinforcement Learning, IRL) [10, 11, 12], directly learning and predicting the actions of the agent using Behavior Cloning Strategies (BC) [13, 14, 15, 16] or hybrid approaches where a discriminator must distinguish between artificial actions from a generator and actual demonstration actions [17, 18]. Diffusion models have recently emerged as state-of-the-art for imitation learning [19, 20, 21, 22, 23]. However, these models often neglect agents’ limited perception, which is critical for accurate imitation. We argue that we need to account for the perception capabilities and observation space of the agents to better model, predict, and imitate agents such as humans. Since actions depend on observations, robots that model limited perception of other agents, improve prediction accuracy.
@@
Figure 2: We employ the diffusion policy learned from human data on a car. Like humans, the car changes its observation direction to look around, detect obstacles, and avoid them while moving toward the goal in real-time and successfully navigates the environment.
@@
We propose that to better model, predict, and learn agents’ motion, it is necessary to correctly model both the observation model of the agent and their policy. There are two key components that affect an agent’s perception of an environment, which ultimately affect an agent’s actions: (a) the field of view and range of the agent’s visual observation, which determines where the agent is looking and what it can potentially perceive, and (b) the likelihood of successfully detecting objects within that observation region by the agent. In this paper, we examine the role of these two components and show that we can infer and estimate these parameters using demonstration trajectories of an agent.
@@
In this work, to better understand the role of an agent’s observation strategy, we create a robotic agent with a limited field of view, view range, and likelihood of observing an obstacle, and use the agent’s motion to infer these quantities. We consider three distinct scenarios, where each scenario involves one unknown: the parameters of the observation space, the probability of detecting obstacles within the agent’s field of view, or the agent’s motion policy. In all scenarios, we show that by accounting for the agent’s inherent observation limitations, we can more accurately model and predict its behavior.
It should be noted that while an end-to-end learning approach could, in theory, capture the complexities of human-like navigation, our modular approach offers several advantages in terms of interpretability, generalization, and robustness. By explicitly modeling the agent’s observation space, detection probability, and motion policy separately, we can better understand decision-making. This structure also enhances generalization, as the learned components can be adapted to different environments without requiring extensive retraining. This decomposition also isolates failures, making them easier to diagnose and mitigate compared to end-to-end learning, where errors can propagate unpredictably.
In all scenarios, we show that by accounting for the agent’s inherent observation limitations, we can more accurately model and predict its behavior. More specifically, we show that to infer the parameters of the observation space, one can use the cross-entropy optimization [24] to estimate the agent’s observation parameters accurately. We then show that using Bayesian optimization [25], the probability of object detection within an agent’s field of view can be easily estimated. Finally, for inferring the policy of the agent, we show that the agent’s policy can be accurately imitated using diffusion models [20]. We illustrate this by training a diffusion policy on human demonstrations of navigating an environment under a restricted field of view. We demonstrate our ability to predict and replicate human behavior in the presence of static and dynamic obstacles. Finally, we deploy the imitated policy to physical hardware, showcasing a robot navigating a cluttered environment and exhibiting human-like behavior by dynamically rotating its field of view to detect obstacles.
@@ II-A Problem Formulation
Consider an agent navigating in an environment with obstacles. The obstacles could be either stationary or moving. We assume that the agent operates under partial observability as it can only perceive a limited portion of the environment at any time. This naturally aligns with a Partially Observable Markov Decision Process (POMDP) framework, where the agent maintains a belief over the true state of the environment based on incomplete observations.
We use $x_{t}\in\mathbb{R}^{n}$ to denote the state of the agent at time $t\in\mathbb{N}\cup\{0\}$, where is $\mathbb{N}$ is the set of natural numbers, and $n$ represents the agent’s state space dimension. Let the control action of the agent at time $t$ be denoted by $u_{t}\in\mathbb{R}^{m}$ where $m$ is the dimension of the control space of the agent. We consider discrete deterministic dynamics for the agent to be denoted by the function $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$, such that:
@@
We assume the robot’s dynamics are known and that the agent has limited observation capabilities. The agent’s observation at time $t$ is denoted via $z_{t}\in\mathbb{R}^{o}$ where $o$ is the dimension of the observation space of the agent. We assume that the agent has an observation space that determines the area of the environment that the agent is looking towards. Furthermore, we assume that the agent has an inherent probability of observing something present in its observation area. With the introduction of this probability, we try to incorporate the scenarios in which humans sometimes miss objects even when they are present in their view. The combined observation space and probability of observation detection determine the agent’s observation at each time, given the agent’s current state.
Let the observation space parameters that determine the agent’s observation area be denoted by $\theta\in\Theta$ where $\Theta$ is the space of observation space parameters. Also, let the probability of observing a particular observation be denoted by $p_{obs}\in[0,1]$. These components define the agent’s POMDP observation model, which we represent as
@@
At each time $t$, based on the observation $z_{t}$, the agent takes a control action $u_{t}$. The probability of taking an action at each time $t$ by the agent is captured through a stochastic policy, which we denote via $\pi:\mathbb{R}^{m}\times\mathbb{R}^{o}\rightarrow[0,1]$,
@@
Running Example: To illustrate this idea, consider an agent with discrete dynamics (1) navigating through an environment toward its desired goal (Fig. 1). The agent perceives the scene with a visual sensor parameterized by the range of vision ($r_{obs}$), and the field of view ($\theta_{obs}$). Obstacles in the field of view are not guaranteed to be noticed (due to lack of visual features), which is captured by the detection probability $p_{obs}$. Furthermore, the orientation in which robot looking, i.e. the orientation of the visual sensor is denoted via $\psi_{obs}$. Note that the motion of the agent in any environment is dependent on the controls inputs chosen by the robot. Also, the agent can choose the orientation of the visual sensor to look in a particular direction. Therefore, we assume that the motion of the agent is governed by a joint policy $\pi(\cdot;\theta_{\pi})$, which controls a) the motion of the robot in the environment, and b) the orientation of the visual sensor denoted by $\psi_{obs}$. With these constraints, the agent follows a POMDP-based navigation strategy, making decisions under uncertainty to minimize collisions while reaching the goal.
Given this example, we look at three key problems: 1) estimating the agent’s observation space parameters, ($r_{obs},\theta_{obs}$), 2) calculating the probability of detecting an obstacle within the field of view, $p_{obs}$, and 3) estimating the agent’s motion policy for both displacement and sensor motion, $\pi(\cdot;\theta_{\pi})$ from a set of agent behavior demonstrations.
To analyze the influence of each component independently, we adopt a modular approach: we estimate each unknown parameter while assuming the other two are known. This structure allows us to isolate the effect of each component on the agent’s behavior and evaluate the accuracy of each estimation method.
@@ II-B Estimating the Observation Space
We assume that the probability of observing an obstacle $p_{obs}$ within an observation space and the agent’s policy parameters $\theta_{\pi}$ are known.
@@
$P(\tau;\theta)=P(x_{0})\Pi_{t=0}^{T-1}P(x_{t+1}|x_{t},u_{t})\int P(z_{t}|x_{t})P(u_{t}|z_{t})dz_{t}$
$=\Pi_{t=0}^{T-1}\int O(z_{t}|x_{t};\theta,p_{obs})\pi(u_{t}|z_{t}(\theta);\theta_{\pi})dz_{t}.$ (4)
From now on, we omit denoting that integration is over $z_{t}$ for the sake of brevity. From (II-B), we can compute the probability of the whole dataset $\mathcal{D}$ as $P(\mathcal{D};\theta)=\sum_{i}^{N}P(\tau_{i};\theta).$ We can alternatively maximize the log-likelihood of the trajectories:
@@
Solving (5) requires taking the gradient of the objective with respect to $\theta$, which would result in having to take the gradient of policy $P(u_{t}|z_{t})$ with respect to $\theta$ where $z_{t}$ depends on $\theta$. This is difficult due to its nested probabilistic dependencies. Furthermore, the presence of an integral over $z_{t}$ makes analytically computing the gradient of the objective intractable. Therefore, we resort to black-box optimization techniques such as Bayesian optimization [25] and the cross-entropy method [24], which find the optimal solutions based on evaluations of the function without requiring us to compute the gradient of the objective. Bayesian optimization is a probabilistic model-based optimization technique used to optimize objective functions, which are expensive to evaluate. The Cross-Entropy Method is another optimization technique that iteratively samples random parameters, evaluates their performance, and refines the parameter distribution toward the optimal region [24]. To estimate the observation space, we use the Cross-Entropy method.
@@ II-C Estimating the Probability of Observation
Under this assumption, we can compute the probability of a trajectory given a value of $p_{obs}$ and find $p_{obs}$ that maximizes the likelihood. Estimating $p_{obs}$ can be cast as the following maximum likelihood estimation problem:
@@
The problem in (6) is similar to that of (5). Therefore, we can use the same black-box function optimization techniques. Since we need to estimate only one parameter $p_{obs}$, we employ Bayesian optimization.
@@ II-D Estimating the Policy of the Agent
Under this assumption, the problem of estimating $\theta_{\pi}$ can be treated as a behavior cloning problem, i.e., we aim to learn the policy $\pi(\cdot;\theta_{\pi})$ parameterized by $\theta_{\pi}$ such that it imitates the provided data set $\mathcal{D}$ in the best possible way. Unlike prior BC approaches that assume full observability, our method conditions the policy on what the agent can perceive, explicitly incorporating limited field of view and probabilistic detection. This leads to a more realistic imitation of the agent behavior. If the policy is a neural network, then $\theta_{\pi}$ can be considered to be the neural network parameters. State of the art in behavior cloning is diffusion policies [21]. These methods model policy generation as a diffusion process, where an initial noisy signal or random trajectory is progressively refined toward optimal actions through iterative steps. As such, we adopt diffusion policies for mimicking the agent’s policy. We formulate policies such as Denoising Diffusion Probabilistic Models (DDPMs) conditioned on the state and observation of the agent. The detailed implementation for this method can be found in [21].
@@ III Numerical Simulations
In this section, we present numerical simulations to showcase the importance of considering agents’ observation and perception capabilities. We consider synthetic settings where we have access to the ground-truth parameters, which enable a comprehensive comparison and evaluation of our method. Then, in the next section, we consider human data and perform hardware experiments that leverage actual human data. We consider a unicycle agent with discrete dynamics:
@@
Figure 3: Schematic figure for the unicycle navigation scenario with partial observations. The bright orange circle represents the unicycle. The gray circles represent the environmental obstacles, and the green represents the goal. The light orange cone represents the observation space of the agent, which is parameterized by the radius of the cone ($r_{obs}$) and the angle of the cone ($\theta_{obs}$). Each obstacle lights up red when the agent observes the obstacle.
@@
First, we consider the problem of estimating observation space parameters of the unicycle agent, $\theta=(r_{obs},\theta_{obs})$. We consider four different scenarios with different $(r_{obs},\theta_{obs})$ values and generate four datasets of 500 trajectories each. During each trajectory generation, we randomly choose the location of the obstacles. We assume that the agent follows a known deterministic policy that is attractor-repulsor in nature, i.e., a policy that guides the agent to move toward the goal while avoiding obstacles. The visualization of one such trajectory is provided in Figure 3. We use the cross-entropy method to solve (5) and estimate the actual parameters $(r_{obs},\theta_{obs})$. For each set of parameters, we provide the ground truth values and the estimated values in Table I. As it can be seen, our method can closely estimate the actual parameter values.
@@
TABLE I: Actual and estimated values of the observation space parameters.
We further examine the trajectories that these estimated $\theta^{*}=(r^{*}_{obs},\theta^{*}_{obs})$ induce to examine how close they are to the demonstration data. We compute the Fréchet distance between the actual and predicted trajectories to investigate their similarity. The Fréchet distance is a good metric for comparing curves because it captures the overall shape and continuity by considering the closest possible continuous matching between the two curves [26]. However, the absolute value of the Fréchet distance can be difficult to interpret in isolation, so we normalize it by the straight-line distance between the start and goal locations to obtain a scale-invariant measure of trajectory similarity.
@@
TABLE II: Average normalized Fréchet distance between 500 actual and predicted trajectories for scenarios in Table I.
The normalized average Fréchet distances between actual and predicted trajectories for all scenarios are provided in Table II. As can be seen, for all four datasets, the estimated parameters are within about $6\%$ of the interval to the original trajectories. These results confirm that by reasoning about the observation space and perception capabilities of the agent, we can mimic the agent behavior and generate agent trajectories that are close to the demonstrated trajectories.
Next, we estimate the probability of observing the obstacles, $p_{obs}$. For this method, we consider the same policy as before, and we further assume access to the actual observation space parameters $(r_{obs},\theta_{obs})$. Same as before, we consider four different scenarios which correspond to four different values of $p_{obs}$ to generate four datasets of 500 trajectories each. Same as before, we randomly choose the locations of obstacles during each trajectory. We employ an open-source Python toolbox for Bayesian optimization [27] to solve (6) and estimate the $p_{obs}$ values. The actual values of $p_{obs}$ and their corresponding estimated values are presented in Table III. Our method can predict the $p_{obs}$ values within about $10\%$ of the actual values.
@@
TABLE III: Actual and estimated values of observation probability $p_{obs}$.
@@
TABLE IV: Average normalized Fréchet distance between 500 actual and predicted trajectories for scenarios in Table III.
Finally, we focus on estimating the agent’s policy. We generate a dataset of 500 trajectories with different obstacle configurations for all trajectories using the aforementioned attractor-repulsor policy with fixed values of $r_{obs},\theta_{obs}$, and $p_{obs}$. We aim to learn a policy that best imitates the dataset. For this purpose, we employ Diffusion policies where we learn an action diffusion policy that takes action based on observation and the agent’s current state. Note that, unlike the previous components where we consider multiple scenarios to estimate different parameters, here, we focus on learning a single policy from the dataset. However, this can easily be extended to different policies. Using the learned diffusion policy, we induce a dataset of predicted trajectories. As before, we compute the normalized Fréchet distances between the actual and predicted trajectories. The average normalized Fréchet distance is $3.47\pm 9.31$, confirming that the resulting trajectories resemble demonstration trajectories.
To further assess the effectiveness and robustness of our modular approach, we evaluate the generalization capability of the learned diffusion policy. Specifically, we train our diffusion policy on a set of observation space parameters and test its performance on previously unseen observation parameters. We then compare the trajectories generated by the learned policy with those produced by the original policy. The results, summarized in Table V, show the average normalized Fréchet distances between the predicted and original trajectories across different observation parameters. Despite variations in the observation space, the learned policy successfully generalizes to new settings, producing trajectories that closely resemble those from the original policy. This demonstrates the adaptability of our modular approach and showcases its ability to capture underlying behavioral patterns while maintaining robust performance across diverse conditions.
@@
TABLE V: Average normalized Fréchet distance between 100 different trajectories generated by the trained diffusion model against the original policy for different observation parameter values.